7.0) Starting tak es about its central is with constant acceleration. In 4.00 (b) the r ecity of the dist
Answer #1

In the problem Magnitude of constant acceleration can be find using below formula

  0. = wot+ -at?

Magnitude of angular velocity at time t can be find using below formula

w=we + at

Given, mist in initially at wat initially angular velocity, w.ro - time, t= 4.00 seconds angular displacement at t o 2.25 rovo = wat + I at 14.13 = (0)14.00) + La (4.00) 14:13 = 0+(1) 14.13 = 8 a = 14113 = 1177 radian second & magnitude of canguler a